OpenStudy (anonymous):

Please check this Find the diameter of the cylinder when:

OpenStudy (anonymous):

Volume=\[12320cm ^{3}\]height=20cm

OpenStudy (anonymous):

im getting \[r^{2}=196\]

OpenStudy (thesecret20111):

Volume of cylinder = pi*r^2*h pi*196 *20 = 12315 which is not quite 12320 but pretty close to it

OpenStudy (thesecret20111):

I'll quickly try working the formula for the volume of cylinder again: Volume of cylinder = pi*r^2*h 12320=pi*r^2*20 Which gives: 616 = pi*r^2 Therefore pi^2 = 196.078889 which is probably what you've gotten. The only problem here is rounding issues. I'd round the number up to 197 you'd probably get a more accurate answer.

OpenStudy (anonymous):

i need the diameter

OpenStudy (anonymous):

@mathslover